Min leaves the house traveling 28 mph in her car. She overtakes Lew in 2 hrs. Lew is riding his bicycle and left the house 5 hours before Min. How fast is Lew traveling
answer is : Lew is traveling 8 mph need a simple break down of how my book got 8mph

Answer by scott8148(6628)

when Min overtakes Lew, they have travelled the same distance __ d=rt

Min drives 28mph for 2h

Lew rides Xmph for (5+2)h

28 * 2 = X * 7 __ 56=7X __ 8=X